# Multihead Finite-State Dimension

### v1

@@ Abstract
We introduce multihead finite-state dimension, a generalization of finite-state dimension in which a group of finite-state agents (the heads) with oblivious, one-way movement rules, each reporting only one symbol at a time, enable their leader to bet on subsequent symbols in an infinite data stream. In aggregate, such a scheme constitutes an $h$-head finite state gambler whose maximum achievable growth rate of capital in this task, quantified using betting strategies called gales, determines the multihead finite-state dimension of the sequence. The 1-head case is equivalent to finite-state dimension as defined by Dai, Lathrop, Lutz and Mayordomo (2004). In our main theorem, we prove a strict hierarchy as the number of heads increases, giving an explicit sequence family that separates, for each positive integer $h$, the earning power of $h$-head finite-state gamblers from that of $(h+1)$-head finite-state gamblers. We prove that multihead finite-state dimension is stable under finite unions but that the corresponding quantity for any fixed number $h>1$ of heads—the $h$-head finite-state predimension—lacks this stability property.
@@ 1 Introduction
This paper introduces multihead finite-state dimension and multihead finite-state strong dimension, two new tools for algorithmic information theory. Our main theorem is a hierarchy result establishing that each additional head strengthens these multihead dimensions.
@@
The present paper was inspired by an ingenious proof by Becher, Carton, and Heiber [4]. Translated into the original gambling (i.e., gale) formulation of finite-state dimension [8], Becher, Carton, and Heiber proved that there is an infinite binary sequence $S$ with the following two properties.
@@
Taken together, items (1) and (2) above imply that equipping a finite-state device with a “probe to the past” as above can make it more powerful than any finite-state device. This is intuitively unsurprising, because any implementation of such a probe involves unbounded memory.
In a line of research over sixty years old [33, 34], a device like $\widehat{G}$ is a two-head finite automaton whose first head is the “leading” (scanning) head of $\widehat{G}$ and whose second head is what we called above a “probe to the past.” The multihead finite automata of this literature (surveyed nicely by Holzer, Kutrib, and Malcher [17]) are typically used to recognize sets of finite strings rather than to compress or gamble on infinite data streams, but the present paper should be regarded as a continuation of this literature as well as a further development of algorithmic dimensions.
Section 3 below defines, for each positive integer $h$, the notion of an $h$-head finite-state gambler ($h$-FSG) in a straightforward way such that a 1-FSG is a finite-state gambler as in [8] and the gambler $\widehat{G}$ above is a 2-FSG. Section 4 then uses section 3 to define, for each positive integer $h$ and each sequence $S\in\Sigma^{\omega}$, the $h$-head finite-state predimension $\dim_{\textup{FS}}^{(h)}(S)$ and the $h$-head finite-state strong predimension $\operatorname{Dim}_{\textup{FS}}^{(h)}(S)$ in such a way that $\dim_{\textup{FS}}^{(1)}(S)=\dim_{\textup{FS}}(S)$ is the finite-state dimension of $S$ as in [8] and $\operatorname{Dim}_{\textup{FS}}^{(1)}(S)=\operatorname{Dim}_{\textup{FS}}(S)$ is the finite-state strong dimension of $S$ as in [3]. It is immediately evident that each of these predimensions is, for fixed $S$, a nonincreasing function of $h$, leading immediately to the definitions of the multihead finite-state dimension
@@
and the multihead finite-state strong dimension
@@
our main theorem says that each of these inequalities is strict for some sequences $S$. This hierarchy theorem is analogous to Yao and Rivest’s nontrivial 1978 proof that, in the context of language recognition, “$k+1$ heads are better than kk” [37], but the proof in our present context is much more involved. For each positive integer $h$, we define a function $F_{h+1}:\Sigma^{\omega}\to\Sigma^{\omega}$, where every sequence $F_{h+1}(S)$ has a structural pattern that is easily exploited by an $(h+1)$-FSG, but each of those $h+1$ heads is essential for this task. We prove that if $R$ is a Martin-Löf random sequence, then no $h$-FSG can gain any significant advantage from the patterns in $F_{h+1}(R)$. Our proof relies on partitioning $F_{h+1}(R)$ into a family of strings of exponentially increasing length, where each string has high conditional Kolmogorov complexity given the information that can be accessed by any (h1)(h-1) heads while the leading head reads that string.
@@
The finite-state dimensions $\dim_{\textup{FS}}$ and $\operatorname{Dim}_{\textup{FS}}$ are stable in the above sense. However, we prove in section 6 that, for each $h\geq 2$, the $h$-head finite-state predimensions $\dim_{\textup{FS}}^{(h)}$ and $\operatorname{Dim}_{\textup{FS}}^{(h)}$ are not stable. We prove in section 7 that the multihead finite-state dimensions $\dim_{\textup{FS}}^{\textup{MH}}$ and $\operatorname{Dim}_{\textup{FS}}^{\textup{MH}}$ are stable. This is why we use the “predimension” terminology.
@@ 2 Preliminaries
  1. 1.
    For $s\in[0,\infty)$, an $s$-gale is a function $d:\Sigma^{*}\rightarrow[0,\infty)$ that satisfies the condition
    $d(w)=|\Sigma|^{-s}\sum_{b\in\Sigma}d(wb)$ (1)
    for all $w\in\Sigma^{*}$.
  2. 2.
    A martingale is a $1$-gale.
@@
    We say that $d$ succeeds on a sequence $S\in\Sigma^{\omega}$ if
    $\limsup_{n\rightarrow\infty}d(S[0..n-1])=\infty.$ (2)
@@
    We say that $d$ succeeds strongly on a sequence $S\in\Sigma^{\omega}$ if
    $\liminf_{n\rightarrow\infty}d(S[0..n-1])=\infty.$ (3)
    The strong success set of $d$ is $S_{\textup{str}}^{\infty}[d]=\{S\in\Sigma^{\omega}\mid d\text{ succeeds strongly on }S\}$.
@@
the minimum length of a program $\pi$ for $\sigma$ given $\tau$ as an auxiliary input. The Kolmogorov complexity of $\sigma$ is
$K(\sigma)=K(\sigma\mid\lambda),$
where $\lambda$ denotes the empty string. Our proofs will only use basic properties of Kolmogorov complexity, some of which we state here. Details and background can be found in [21].
  • For all $\tau\in\{0,1\}^{*}$,
    $K(\tau)\leq|\tau|+2\log|\tau|+O(1).$ (4)
  • For every computable function $f$ and all $\sigma,\tau\in\{0,1\}^{*}$,
@@
    This property is known as the symmetry of information.
A sequence $R\in\mathbf{C}$ is Martin-Löf random if there is some constant $c\in\mathbb{N}$ such that, for all $n\in\mathbb{N}$, we have $K(R[0..n-1])\geq n-c$.
3 Multihead Finite-State Gamblers
We now define multihead finite-state gamblers, which generalize 1-account finite-state gamblers as defined in [8].
@@ 3 Multihead Finite-State Gamblers
Let $h\in\mathbb{Z}^{+}$. An $h$-head finite-state gambler ($h$-FSG) is a 7-tuple $G=(T\times Q,\Sigma,\delta,\mu,\beta,(t_{0},q_{0}),c_{0})$, where
@@
Intuitively, we regard $G$ as having $h$ heads: a leading head which advances through the input sequence in discrete steps, and h1h-1 trailing heads, some subset of which advance in each step. The positions of the heads are governed by the positional states—the members of TT—and the function $\mu$, which dictates, in each positional state, which of the trailing heads advance. More formally, for $r\in\mathbb{N}$, let $\delta_{T}^{r}:T\to T$ denote $r$ iterated applications of $\delta_{T}$. Then the position vector of $G$ after $n$ steps, which tells us the positions of all trailing heads, is determined by the function $\pi:\mathbb{N}\to\mathbb{N}^{h-1}$ defined recursively by $\pi(0)=(0,\ldots,0)$ and, for all $n\in\mathbb{N}$,
$\pi(n+1)=\pi(n)+\mu\left(\delta_{T}^{n}(t_{0})\right).$
We emphasize that because there are infinitely many positional vectors, this is not a finite-state computational model; rather, it generalizes a finite-state model.
For $n\in\mathbb{N}$ and $S\in\Sigma^{\omega}$, let $(\pi_{1},\ldots,\pi_{h-1})=\pi(n)$ and
$\vec{\sigma}_{n}=(S[\pi_{1}],\ldots,S[\pi_{h-1}],S[n]).$
Then the state sequence $(t_{0},q_{0}),(t_{1},q_{1}),(t_{2},q_{2}),\ldots$ of $G$ on input sequence $S\in\Sigma^{\omega}$ is given, for all $n\in\mathbb{N}$, by
@@
Note that the head movements are oblivious in the sense of being data-independent, which, together with the finite-state condition, implies that they each move at an essentially constant rate, in the following sense.
@@
Let $D$ be the set of positional states that appeared before entering the cycle. Note that $D$ and $L$ are disjoint. After $n$ steps, the cycle has been repeated (n|D|)/|L|(n-|D|)/|L| times. For each trailing head $i$, let $\sigma_{i}\in[0,1]$ be such that head $i$ advances $\sigma_{i}|L|$ times during the cycle. Then the position of head $i$ after $n$ steps is bounded below by
$\sigma_{i}|L|\lfloor(n-|D|)/|L|\rfloor\geq\sigma_{i}n-|D|-|L|\geq\sigma_{i}n-|T|$
and above by
$|D|+\sigma_{i}|L|\lceil(n-|D|)/|L|\rceil\leq\sigma_{i}n+|D|+|L|\leq\sigma_{i}n+|T|.\qed$
4 Multihead Finite-State Dimensions
@@
The martingale of an $h$-FSG $G$ is the function $d_{G}:\Sigma^{*}\rightarrow[0,\infty)$ defined recursively by $d_{G}(\lambda)=c_{0}$ and
$d_{G}(wb)=|\Sigma|d_{G}(w)\beta(q_{|w|})(b),$
where $q_{|w|}$ is defined as in (7).
@@
For $s\in[0,\infty)$, the $s$-gale of an $h$-FSG $G$ is the function $d_{G}^{(s)}:\Sigma^{*}\rightarrow[0,\infty)$ defined, for all $w\in\Sigma^{*}$, by
$d_{G}^{(s)}(w)=|\Sigma|^{(s-1)|w|}d_{G}(w).$
@@
    The $h$-head finite-state predimension of $X$ is
    $\dim_{\textup{FS}}^{(h)}(X)=\inf\mathcal{G}_{h}(X).$
@@
    The $h$-head finite-state strong predimension of $X$ is
    $\operatorname{Dim}_{\textup{FS}}^{(h)}(X)=\inf\mathcal{G}_{h}^{\textup{str}}(X).$
@@
    The multihead finite-state dimension of $X$ is
    $\dim_{\textup{FS}}^{\textup{MH}}(X)=\inf_{h\in\mathbb{Z}^{+}}\dim_{\textup{FS}}^{(h)}(X).$
@@
    The multihead finite-state strong dimension of $X$ is
    $\operatorname{Dim}_{\textup{FS}}^{\textup{MH}}(X)=\inf_{h\in\mathbb{Z}^{+}}\operatorname{Dim}_{\textup{FS}}^{(h)}(X).$
Both $h$-head finite-state predimension and multihead finite-state dimension are monotone, in the following sense.
@@
The $h$-head finite-state predimension, $h$-head finite-state strong predimension, multihead finite-state dimension, and multihead finite-state strong dimension of an individual sequence $S\in\Sigma^{\omega}$ are the $h$-head finite-state predimension, $h$-head finite-state strong predimension, multihead finite-state dimension, and multihead finite-state strong dimension, respectively, of the singleton $\{S\}$.

### v2

# Multi-Head Finite-State Dimension

@@ Abstract
We introduce multi-head finite-state dimension, a generalization of finite-state dimension in which a group of finite-state agents (the heads) with oblivious, one-way movement rules, each reporting only one symbol at a time, enable their leader to bet on subsequent symbols in an infinite data stream. In aggregate, such a scheme constitutes an $h$-head finite state gambler whose maximum achievable growth rate of capital in this task, quantified using betting strategies called gales, determines the multi-head finite-state dimension of the sequence. The 1-head case is equivalent to finite-state dimension as defined by Dai, Lathrop, Lutz and Mayordomo (2004). In our main theorem, we prove a strict hierarchy as the number of heads increases, giving an explicit sequence family that separates, for each positive integer $h$, the earning power of $h$-head finite-state gamblers from that of $(h+1)$-head finite-state gamblers. We prove that multi-head finite-state dimension is stable under finite unions but that the corresponding quantity for any fixed number $h>1$ of heads—the $h$-head finite-state predimension—lacks this stability property.
@@ 1 Introduction
This paper introduces multi-head finite-state dimension and multi-head finite-state strong dimension, two new tools for algorithmic information theory. Our main theorem is a hierarchy result establishing that each additional head strengthens these multi-head dimensions.
@@
The present paper was motivated by a long-standing open question and inspired by an ingenious proof. The long-standing open question is whether the finite-state setting is the only level of effectivity in which randomness is equivalent to full dimension. Bourke, Hitchcock, and Vinodchandran [5] noted that the equivalence of finite-state randomness and finite-state full dimension follows from a theorem of Schnorr and Stimm [42]. In contrast, there is no such equivalence for algorithmic randomness and dimension, computable randomness and dimension, or resource-bounded randomness and dimension; in all these settings, there exist non-random sequences with full dimension.
The search for other settings where this equivalence holds has focused on notions of randomness and dimension that are “highly effective,” defined using gamblers that are computationally weaker than polynomial-time algorithms. Examples include the settings of pushdown gamblers, Lempel-Ziv gamblers, and perhaps logspace gamblers [1, 2, 10, 26, 38, 32], but it is not yet known whether randomness and full dimension are equivalent in those settings. Here our objective is to investigate a setting that is as close as possible to, but not quite as effective as, the setting of finite-state gamblers.
The ingenious proof that inspired our choice of a setting to investigate is due to Becher, Carton, and Heiber [4]. Translated into the original gambling (i.e., gale) formulation of finite-state dimension [8], Becher, Carton, and Heiber proved that there is an infinite binary sequence $S$ with the following two properties.
@@
Taken together, items (1) and (2) above imply that equipping a finite-state gambler with a “probe to the past” as above can make it more powerful than any finite-state gambler. This is intuitively unsurprising, because implementing such a probe requires access to an unbounded amount of past data.
In a line of research over sixty years old [40, 41], a device like $\widehat{G}$ is a two-head finite automaton whose first head is the “leading” (scanning) head of $\widehat{G}$ and whose second head is what we called above a “probe to the past.” The multi-head finite automata of this literature (surveyed nicely by Holzer, Kutrib, and Malcher [21]) are typically used to recognize sets of finite strings rather than to compress or gamble on infinite data streams, but the present paper should be regarded as a continuation of this literature as well as a further development of algorithmic dimensions.
As noted above, the goal here is to investigate a setting that is as close as possible to, but not quite as effective as, the setting of finite-state gamblers. Accordingly, we require the trailing heads in our multi-head finite-state gamblers to be as weak as possible: They can only move forward (or hold still), and they are oblivious in the sense that their movements do not depend on the data that they are scanning. (We are using the terminology of oblivious Turing machines here [39]. In the terminology of [21], our trailing heads are data-independent.) We trust that other, less restrictive variants of our multi-head model will be investigated in due time.
Section 3 below defines, for each positive integer $h$, the notion of an $h$-head finite-state gambler ($h$-FSG) in a straightforward way such that a 1-FSG is a finite-state gambler as in [8] and the gambler $\widehat{G}$ above is a 2-FSG. Section 4 then uses section 3 to define, for each positive integer $h$ and each sequence $S\in\Sigma^{\omega}$, the $h$-head finite-state predimension $\dim_{\textup{FS}}^{(h)}(S)$ and the $h$-head finite-state strong predimension $\operatorname{Dim}_{\textup{FS}}^{(h)}(S)$ in such a way that $\dim_{\textup{FS}}^{(1)}(S)=\dim_{\textup{FS}}(S)$ is the finite-state dimension of $S$ as in [8] and $\operatorname{Dim}_{\textup{FS}}^{(1)}(S)=\operatorname{Dim}_{\textup{FS}}(S)$ is the finite-state strong dimension of $S$ as in [3]. It is immediately evident that each of these predimensions is, for fixed $S$, a nonincreasing function of $h$, leading immediately to the definitions of the multi-head finite-state dimension
@@
and the multi-head finite-state strong dimension
@@
We note that our model is definitionally robust, in that N. Lutz [31] has very recently shown that our multi-head finite-state predimensions and dimensions are equivalent to natural notions of multi-head finite-state compressibility.
@@
our main theorem says that each of these inequalities is strict for some sequences $S$.
This hierarchy theorem is analogous to Yao and Rivest’s nontrivial 1978 proof that, in the context of language recognition, “$k+1$ heads are better than kk” [45] and the subsequent line of hierarchy theorems for the recognition power of multi-head finite automata in two-way [33, 20] or data-independent [20, 19, 12, 13] settings.
For each positive integer $h$, we define a function $F_{h+1}:\Sigma^{\omega}\to\Sigma^{\omega}$, where every sequence $F_{h+1}(S)$ has a structural pattern that is easily exploited by an $(h+1)$-FSG, but each of those $h+1$ heads is essential for this task. We prove that if $R$ is a Martin-Löf random sequence, then no $h$-FSG can gain any significant advantage from the patterns in $F_{h+1}(R)$. Our proof relies on partitioning $F_{h+1}(R)$ into a family of strings of exponentially increasing length, where each string has high conditional Kolmogorov complexity given the information that can be accessed by any (h1)(h-1) heads while the leading head reads that string.
@@
The finite-state dimensions $\dim_{\textup{FS}}$ and $\operatorname{Dim}_{\textup{FS}}$ are stable in the above sense. However, we prove in section 6 that, for each $h\geq 2$, the $h$-head finite-state predimensions $\dim_{\textup{FS}}^{(h)}$ and $\operatorname{Dim}_{\textup{FS}}^{(h)}$ are not stable. We prove in section 7 that the multi-head finite-state dimensions $\dim_{\textup{FS}}^{\textup{MH}}$ and $\operatorname{Dim}_{\textup{FS}}^{\textup{MH}}$ are stable. This is why we use the “predimension” terminology.
@@ 2 Preliminaries
Following the standard template for effectivizations of fractal dimensions [27, 28, 3], we will base our dimension notions on the success of functions called gales.
@@
For $s\in[0,\infty)$, an $s$-gale is a function $d:\Sigma^{*}\to[0,\infty)$ that satisfies the condition
$d(w)=|\Sigma|^{-s}\sum_{b\in\Sigma}d(wb)$
for all $w\in\Sigma^{*}$. A martingale is a $1$-gale.
Intuitively, the value of a gale is interpreted as the capital of a gambler betting on successive symbols in a sequence. The number $s$ parameterizes the favorability of the betting environment. A martingale, where $d(w)$ is the average, over all symbols $b$, of $d(wb)$, corresponds to a perfectly fair betting environment, in which a maximally conservative gambler who places equal bets on all possible one-symbol extensions of $w$ can maintain constant capital. Parameters $s<1$ correspond to less favorable betting environments where “the house takes a cut.” In these environments, the gambler would need to exploit some knowledge about the sequence—or be lucky—to maintain constant capital. For a gale to succeed, though, we require that the gambler’s capital is unbounded, not merely constant.
@@
    We say that $d$ succeeds strongly on a sequence $S\in\Sigma^{\omega}$ if $\liminf_{n\rightarrow\infty}d(S[0..n-1])=\infty$. The strong success set of $d$ is $S_{\textup{str}}^{\infty}[d]=\{S\in\Sigma^{\omega}\mid d\text{ succeeds strongly on }S\}$.
In general, succeeding on a sequence becomes more difficult as the sequence becomes more complex, as the computational resources available to the gale become more restricted, and as the value $s$ decreases.
@@
While the success of gales quantifies the predictability of sequences, Kolmogorov complexity quantifies the compressibility of strings. These are closely related concepts, and several of our proofs rely on bounding the compressibility of parts of a sequence to derive bounds on the sequence’s overall predictability. Our proofs will use the basic properties of Kolmogorov complexity stated below. Further details and background on Kolmogorov complexity can be found in [25].
@@
that is, the minimum length of a program $\pi$ for $\sigma$ given $\tau$ as an auxiliary input. The Kolmogorov complexity of $\sigma$ is $K(\sigma)=K(\sigma\mid\lambda)$, where $\lambda$ denotes the empty string.
  • Computable functions do not add information: For every computable function $f$ and all $\sigma,\tau\in\{0,1\}^{*}$,
@@
A sequence $R\in\mathbf{C}$ is Martin-Löf random if there is some constant $c\in\mathbb{N}$ such that, for all $n\in\mathbb{N}$, we have $K(R[0..n-1])\geq n-c$; in this sense, such sequences are incompressible. We will construct sequences based on Martin-Löf random sequences, and our constructed sequences will inherit some of this incompressibility.
3 Multi-Head Finite-State Gamblers
We now define multi-head finite-state gamblers. These are a generalization of the 1-account finite-state gamblers defined in [8]; our definition here is self-contained. We first define the formal components of a gambler, then describe its movements and operation.
The gambler’s state space will be the Cartesian product of two sets, $T$ and $Q$. The $T$ component of the state will govern the heads’ movements, while the $Q$ component will govern its betting behavior. We will place a requirement on the transition function $\delta$ that it can be decomposed into separate transition functions $\delta_{T}$ and $\delta_{Q}$ that act on $T$ and $Q$, respectively, where $\delta_{T}$ does not depend on the symbols read. This has the effect of isolating the $T$ component of the gambler’s state from the tape contents, which makes the head movements data-independent.
@@ 3 Multi-Head Finite-State Gamblers
Let $h\in\mathbb{Z}^{+}$. An $h$-head finite-state gambler ($h$-FSG) is a 7-tuple
$G=(T\times Q,\Sigma,\delta,\mu,\beta,(t_{0},q_{0}),c_{0}),$
where
@@
Intuitively, we regard $G$ as having $h$ heads: a leading head which advances through the input sequence in discrete steps, and h1h-1 trailing heads, some subset of which advance in each step. The positions of the heads are governed by the positional states—the members of TT—and the function $\mu$, which dictates, in each positional state, which of the trailing heads advance. More formally, for $r\in\mathbb{N}$, let $\delta_{T}^{r}:T\to T$ denote $r$ iterated applications of $\delta_{T}$. Then the positional vector of $G$ after $n$ steps, which tells us the positions of all trailing heads, is determined by the function $\pi:\mathbb{N}\to\mathbb{N}^{h-1}$ defined by $\pi(0)=(0,\ldots,0)$ and, for all $n\in\mathbb{N}$, recursively by $\pi(n+1)=\pi(n)+\mu\left(\delta_{T}^{n}(t_{0})\right)$.
For $n\in\mathbb{N}$ and $S\in\Sigma^{\omega}$, let $(\pi_{1},\ldots,\pi_{h-1})=\pi(n)$ and $\vec{\sigma}_{n}=(S[\pi_{1}],\ldots,S[\pi_{h-1}],S[n])$. Then the state sequence $(t_{0},q_{0}),(t_{1},q_{1}),(t_{2},q_{2}),\ldots$ of $G$ on input sequence $S\in\Sigma^{\omega}$ is given, for all $n\in\mathbb{N}$, by
@@
Note that the head movements are oblivious in the sense of being data-independent, which, together with the finite-state condition, implies that each of them moves at an essentially constant rate, in the following sense.
@@
Let $D$ be the set of positional states that appeared before entering the cycle. Note that $D$ and $L$ are disjoint. After $n$ steps, the cycle has been repeated (n|D|)/|L|(n-|D|)/|L| times. For each trailing head $i$, let $\sigma_{i}\in[0,1]$ be such that head $i$ advances $\sigma_{i}|L|$ times during the cycle. Then the position of head $i$ after $n$ steps is bounded below by $\sigma_{i}|L|\left\lfloor\frac{n-|D|}{|L|}\right\rfloor\geq\sigma_{i}n-|D|-|L|\geq\sigma_{i}n-|T|$ and above by $|D|+\sigma_{i}|L|\left\lceil\frac{n-|D|}{|L|}\right\rceil\leq\sigma_{i}n+|D|+|L|\leq\sigma_{i}n+|T|$. ∎
4 Multi-Head Finite-State Dimensions
We now define $h$-head finite-state (strong) predimension—which generalize the finite-state dimension of [8] and the finite-state strong dimension of [3]—and multi-head finite-state (strong) dimension. Definitions 4.14.4 continue to follow the standard template for effectivizing fractal dimensions [27, 28, 3], differing from the corresponding definitions in [8, 3] only in that the gamblers now have $h$ heads.
Briefly, the approach is to define the martingale and $s$-gales induced by the betting function of a specific gambler, then consider the range of parameters $s$ for which it is possible for a gambler’s $s$-gale to (strongly) succeed on a given set. Recalling that $s$ parameterizes the favorability of the betting environment, this means we are quantifying how unfavorable the betting environment can become before it becomes impossible for an $h$-FSG to succeed on the set. Intuitively, a gambler can succeed in a more unfavorable environment if the set only contains simple, predictable sequences, so such sets will tend to have lower predimension or dimension.
@@ 4 Multi-Head Finite-State Dimensions
The martingale of an $h$-FSG $G$ is the function $d_{G}:\Sigma^{*}\rightarrow[0,\infty)$ defined recursively by $d_{G}(\lambda)=c_{0}$ and $d_{G}(wb)=|\Sigma|d_{G}(w)\beta(q_{|w|})(b)$, where $q_{|w|}$ is defined as in (3).
By rescaling this martingale, we derive corresponding $s$-gales.
@@
For $s\in[0,\infty)$, the $s$-gale of an $h$-FSG $G$ is the function $d_{G}^{(s)}:\Sigma^{*}\rightarrow[0,\infty)$ defined, for all $w\in\Sigma^{*}$, by $d_{G}^{(s)}(w)=|\Sigma|^{(s-1)|w|}d_{G}(w)$.
Given a set of sequences, we may then consider the range of parameters $s$ for which $h$-FSGs with (strongly) successful $s$-gales exist.
@@
The infima of these ranges then give us our predimensions; the $h$-head finite-state (strong) predimension of a set $X$ is the infimum of the parameters $s$ such that an $h$-FSG can have an $s$-gale that (strongly) succeeds on $X$.
@@
    The $h$-head finite-state predimension of $X$ is $\dim_{\textup{FS}}^{(h)}(X)=\inf\mathcal{G}_{h}(X)$.
@@
    The $h$-head finite-state strong predimension of $X$ is $\operatorname{Dim}_{\textup{FS}}^{(h)}(X)=\inf\mathcal{G}_{h}^{\textup{str}}(X)$.
@@
    The multi-head finite-state dimension of $X$ is $\dim_{\textup{FS}}^{\textup{MH}}(X)=\inf_{h\in\mathbb{Z}^{+}}\dim_{\textup{FS}}^{(h)}(X)$.
@@
    The multi-head finite-state strong dimension of $X$ is $\operatorname{Dim}_{\textup{FS}}^{\textup{MH}}(X)=\inf_{h\in\mathbb{Z}^{+}}\operatorname{Dim}_{\textup{FS}}^{(h)}(X)$.
Both $h$-head finite-state predimension and multi-head finite-state dimension are monotone, in the following sense.
@@
The $h$-head finite-state predimension, $h$-head finite-state strong predimension, multi-head finite-state dimension, and multi-head finite-state strong dimension of an individual sequence $S\in\Sigma^{\omega}$ are the $h$-head finite-state predimension, $h$-head finite-state strong predimension, multi-head finite-state dimension, and multi-head finite-state strong dimension, respectively, of the singleton $\{S\}$.